OpenStudy (anonymous):

The Greens Manufacturing Company makes 60% of a particular type of camera lens, the Parsons Company makes 15% of them, and the Ratten Company makes the remaining 25%. Of all the camera lenses, 5% are made by Greens and are defective, 10% are made by Parsons and are defective, and 6% are made by Ratten and are defective. Lens Defective Not Defective Total Greens 5% 55% 60% Parsons 10% 5% 15% Ratten 6% 19% 25% Total

OpenStudy (anonymous):

If a camera lens is randomly selected from the general population of all lenses, what is the probability that it was NOT made by the Ratten Company, given that it is defective? 0.22 0.36 0.49

OpenStudy (anonymous):

P(made by ratten) = 1/4 = 25%

OpenStudy (anonymous):

P(not madde by ratten) = 3/4

OpenStudy (anonymous):

Greens = 60% Parsons = 15% Ratten = 25 % ------------- Total = 100%

OpenStudy (anonymous):

Defective Green = 5% Defective Parsons = 10% ------------------------ total defective: 15%

OpenStudy (anonymous):

55% of of Greens fine 5 % of Parsons fine 19% of Ratten fine I think you have already given answer in your question